Exercise 9: A researcher is interested in the relationship between oral contraceptive (OC) use and blood pressure in women. SType or paste question hereresearcher assumes that the two groups are independent and normal will equal variances. 1. Using the p-value, does the underl
Answer #1

1. The hypothesis being tested is:

H0: µ1 = µ2

H1: µ1 ≠ µ2

sp2 = (10 - 1)*15.34^2 + (21 - 1)*18.23^2/(10 + 21 - 2) = 302.22408

se = \sqrt{} 302.22408*(1/10 + 1/21) = 6.67937

The test statistic, t = (132.86 - 127.44)/6.67937 = 0.811

The p-value is 0.4237.

Since the p-value (0.4237) is greater than the significance level (0.05), we fail to reject the null hypothesis.

Therefore, we cannot conclude that µ1 ≠ µ2.

2. The 95% confidence interval will be:

= (132.86 - 127.44) ± 2.05* 6.67937

= (-8.24, 19.08)